Rank the elements in Sr(ClO3)2 from highest to lowest % by mass.
Answer #1

The elements present in the compound Sr(ClO3)2 or SrCl2O6 are:

Sr: atomic mass = 87.62 amu , # of atoms present = 1, contribution to mass of compound = 87.62 amu

Cl: atomic mass = 35.45 amu , # of atoms present = 2, contribution to mass of compound = 2x 35.45 amu = 70.9 amu

O : atomic mass = 16.0 amu , # of atoms present = 6, contribution to mass of compound = 6 x16 amu = 96 amu

mass of the compound = (87.62 + 70.9 + 96 ) amu = 254.52 amu

Clearly mass % of O is highest, followed by Sr and then Cl.

mass % of O = ( 96 / 254.52 ) x 100 = 37.72 %

mass % of Sr= ( 87.62 / 254.52 ) x 100 = 34.42 %

mass % of Cl = (70.9 / 254.52 ) x 100 = 27.86 %

*kindly submit feedback